# Rigidity Hypothesis for the Structural Asymmetry of the Bounded Solutions of a Parametric Complex Linear Differential Equation

### v6

@@ Abstract
In this manuscript, we introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter lying in the critical strip. We identify a rigidity hypothesis along the critical line, which establishes a structural asymmetry: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$. This rigidity hypothesis is satisfied when the non-homogeneous term is the fractional part function.
@@ 3 Application to the $\zeta$-Function
The function $\eta(t)=\{t\}$, the fractional part of $t$, is a locally integrable function and satisfies $0\leq\eta(t)<1$. We will see in the following proposition that the hypothesis (H) is also satisfied
Proposition 7.
@@
When $\eta(t)=\{t\}$ the rigidity hypothesis (H) is satisfied.
Proof.
Given the second item of Remark 6, we prove that : For all $\beta\in\mathbb{R}$ and $t\geq 1$ we have $\Re\big(f_{\beta}(t)\big)\geq 0$, where
$f_{\beta}(t):=\ln(t)-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\,du.$
We have
$f_{\beta}(t)=-\int_{1}^{t}u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du+\int_{1}^{t}u^{-1}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\Big]\,du,$
equivalent to
$f_{\beta}(t)=\int_{1}^{t}u^{-\frac{1}{2}}h_{\beta}(u)\,du=t^{-\frac{1}{2}}\int_{1}^{t}h_{\beta}(u)\,du+\frac{1}{2}\int_{1}^{t}u^{-\frac{3}{2}}\int_{1}^{u}h_{\beta}(v)\,dv\,du,$
where
$h_{\beta}(u):=-u^{-1-i\,\beta}\,\{u\}+u^{-\frac{1}{2}}\Big[1+(\frac{1}{2}-i\beta)\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\Big].$
Then it is sufficient to prove that $\int_{1}^{t}\Re\big(h_{\beta}(u)\big)\,du\geq 0$ for all $t\geq 1$. We have
$\int_{1}^{t}h_{\beta}(u)\,du$ $=2t^{\frac{1}{2}}-2-\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du$
$+(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-\frac{1}{2}}\int_{1}^{u}v^{-\frac{3}{2}-i\,\beta}\,\{v\}\,dv\,du$
$=2t^{\frac{1}{2}}-2-\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du$
$+2(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du-2(\frac{1}{2}-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du,$
We obtained
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\{u\}\,du$
$-(1-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\{u\}\,du.$
Since $\{u\}=u-\lfloor u\rfloor$, and since
$(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{1}{2}-i\,\beta}\,du-(1-i\beta)\int_{1}^{t}u^{-i\,\beta}\,\,du=0,$
then
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\int_{1}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\lfloor u\rfloor\,du$
$+(1-i\beta)\int_{1}^{t}u^{-1-i\,\beta}\,\lfloor u\rfloor\,du,$
which can be written as
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1$
$-(\frac{1}{2}-i\beta)t^{\frac{1}{2}}\Big[\lfloor t\rfloor\int_{\lfloor t\rfloor}^{t}\,u^{-\frac{3}{2}-i\,\beta}\,\,du+\sum_{n=1}^{\lfloor t\rfloor-1}n\int_{n}^{n+1}\,u^{-\frac{3}{2}-i\,\beta}\,du\Big]$
$+(1-i\beta)\Big[\lfloor t\rfloor\int_{\lfloor t\rfloor}^{t}u^{-1-i\,\beta}\ du+\sum_{n=1}^{\lfloor t\rfloor-1}n\int_{n}^{n+1}\,u^{-1-i\,\beta}\,du\Big].$
Integrate, to obtain
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-t^{\frac{1}{2}}\frac{\frac{1}{2}-i\beta}{-\frac{1}{2}-i\,\beta}\lfloor t\rfloor\big(t^{-\frac{1}{2}-i\,\beta}-\lfloor t\rfloor^{-\frac{1}{2}-i\,\beta}\big)$
$+\frac{1-i\beta}{-i\,\beta}\lfloor t\rfloor\big(t^{-i\,\beta}-\lfloor t\rfloor^{-i\,\beta}\big)$
$-\frac{\frac{1}{2}-i\beta}{-\frac{1}{2}-i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor-1}n\big((n+1)^{-\frac{1}{2}-i\,\beta}-n^{-\frac{1}{2}-i\,\beta}\big)$
$+\frac{1-i\beta}{-i\,\beta}\sum_{n=1}^{\lfloor t\rfloor-1}n\big((n+1)^{-i\,\beta}-n^{-i\,\beta}\big).$
Using telescopic series, one obtains
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}-\frac{1-i\beta}{i\,\beta}\Big)\lfloor t\rfloor t^{-i\,\beta}$
$-\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\Big(1+\sum_{n=1}^{\lfloor t\rfloor-1}(n+1)^{-\frac{1}{2}-i\,\beta}\Big)$
$+\frac{1-i\beta}{i\,\beta}\Big(1+\sum_{n=1}^{\lfloor t\rfloor-1}(n+1)^{-i\,\beta}\Big),$
which is equivalent to,
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1+\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}-\frac{1-i\beta}{i\,\beta}\Big)\lfloor t\rfloor t^{-i\,\beta}$
$-\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}+\frac{1-i\beta}{i\,\beta}\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}.$
or,
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)\{t\}t^{-i\,\beta}$
$+t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}-\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}$
$+\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)t^{1-i\,\beta}$
$-\frac{1}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}+\frac{1}{i\,\beta}\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}.$
Use the Abel summation formula to the last line, we obtain
$\frac{1}{2}\int_{1}^{t}h_{\beta}(u)\,du$ $=t^{\frac{1}{2}}-1-\Big(\frac{1}{\frac{1}{2}+i\,\beta}-\frac{1}{i\,\beta}\Big)\{t\}t^{-i\,\beta}$
$+t^{\frac{1}{2}}\sum_{n=1}^{\lfloor t\rfloor}n^{-\frac{1}{2}-i\,\beta}-\sum_{n=1}^{\lfloor t\rfloor}n^{-i\,\beta}$
$+\frac{1}{\frac{1}{2}+i\,\beta}t^{\frac{1}{2}}+t^{\frac{1}{2}}\int_{1}^{t}u^{-\frac{3}{2}-i\beta}\{u\}\,du$
$-\frac{1}{1-i\,\beta}-\int_{1}^{t}u^{-1-i\beta}\{u\}\,du.$
A direct estimate show that $\Re\big(f_{\beta}(t)\big)>0$, for all $t>1$. ∎
Then the rigidity hypothesis (H) is satisfied. By definition of the function $\mu$ in the Definition 2, we have
@@
$\frac{1-s}{s}\zeta(s):=-1+\mu(s),\quad\frac{\bar{s}}{1-\bar{s}}\zeta(1-\bar{s}):=-1+\mu(1-\bar{s}).$ (9)
Since the hypothesis (H) is satisfied, by the Theorem 5 when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$. The non-trivial zeros of the Riemann zeta function must lie on the critical line. The results provides an answer to the Dynamical Conjecture formulated in [2].

### v7

@@ Abstract
In this manuscript, we introduce a family of parametrized non-homogeneous linear complex differential equations on $1,\infty)$, depending on a complex parameter lying in the critical strip. We identify a rigidity hypothesis along the critical line, which establishes a structural asymmetry: if two solutions with the same initial condition equal to 11, corresponding respectively to the parameters $s$ and $1-\bar{s}$, are both bounded on $1,+\infty)$, then $\Re(s)=\frac{1}{2}$. This provides a theoretical framework when the non-homogeneous term is the fractional part function.
@@ 3 Application to the $\zeta$-Function
The function $\eta(t)=\{t\}$, the fractional part of $t$, is a locally integrable function and satisfies $0\leq\eta(t)<1$. The following arguments suggests that the rigidity hypothesis (H) is satisfied. The exact calculation is framed within the context of number theory and the study of the discrete zeta function along the critical line.
1. Asymptotic Invariance under Critical Oscillation ($\beta\to+\infty$):
By a application of the Riemann-Lebesgue Lemma, the oscillatory integral terms vanish relative to the logarithmic term as $\beta$ increases. It follows that:
$\lim_{\beta\to+\infty}\mathfrak{R}(f_{\beta}(t))=\ln(t).$
@@ 1. Asymptotic Invariance under Critical Oscillation ($\beta\to+\infty$):
Therefore, for each fixed $t>1$, the rigidity hypothesis (H) is asymptotically satisfied when $\beta\to+\infty$.
2. Dominance of the Zeta Rigidity Term: The function $f_{\beta}$ can be written as
$f_{\beta}(t)$ $=\ln(t)\Big(1+\big(\frac{1}{2}-i\beta\big)\int_{1}^{+\infty}u^{-\frac{3}{2}-i\beta}\{u\}du\Big)$
$-\big(\frac{1}{2}-i\beta\big)\ln(t)\int_{t}^{+\infty}u^{-\frac{3}{2}-i\beta}\{u\}du$
$-\int_{1}^{t}u^{-\frac{3}{2}-i\beta}\{u\}du-\big(\frac{1}{2}-i\beta\big)\int_{1}^{t}\ln(u)u^{-\frac{3}{2}-i\beta}\{u\}\,du.$
Considering the representation of the Riemann zeta function $\zeta$ defined by Abel’s summation formula [[1], page 14, Equation 2.1.5], we have
$\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\beta}\zeta(\frac{1}{2}+i\beta)=1+\big(\frac{1}{2}-i\beta\big)\int_{1}^{+\infty}u^{-\frac{3}{2}-i\beta}\{u\}du,$
Hence, for large $t$, the dynamics of $\Re\big(f_{\beta}(t)\big)$ are governed by
$\Re\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\beta}\zeta(\frac{1}{2}+i\beta)\Big)\ln(t)$
Numerical investigation of the mapping $g(\beta)=-\Re\Big(\frac{\frac{1}{2}-i\beta}{\frac{1}{2}+i\beta}\zeta(\frac{1}{2}+i\beta)\Big)$ reveals two fundamental properties:
  • $g(\beta)\geq 0$ for all $\beta\in\mathbb{R}$.
  • $g(\beta)=0$ if and only if $s=\frac{1}{2}+i\beta$ is a non-trivial zero of the Riemann zeta function.
By definition of the function $\mu$ in the Definition 2, we have
@@ 2. Dominance of the Zeta Rigidity Term: The function $f_{\beta}$ can be written as
$\frac{1-s}{s}\zeta(s):=-1+\mu(s),\quad\frac{\bar{s}}{1-\bar{s}}\zeta(1-\bar{s}):=-1+\mu(1-\bar{s}).$
The hypothesis (H) and the Theorem 5 implies that when $s\in B$ with $\Re(s)\neq\frac{1}{2}$ we get: $\big(\mu(s),\ \mu(1-\bar{s})\big)\neq(1,1)$. This suggests that the non-trivial zeros of the Riemann zeta function must lie on the critical line, providing an answer to the Dynamical Conjecture formulated in [2].